OpenStudy (anonymous):

Evaluate the function at the indicated values. 1.) g(x)=3x^2-4 a.g(x^2) b.g(3x^2-4) c.g(x-h)

OpenStudy (joannablackwelder):

To evaluate the function at the given values, you would substitute whatever is in the parentheses for x in the function and simplify.

OpenStudy (anonymous):

can you give my an example?

OpenStudy (joannablackwelder):

Sure. g(x^3)=3(x^3)^2-4 which simplifies to 3x^6-4

OpenStudy (anonymous):

let me try :D

OpenStudy (joannablackwelder):

Or g(x+2)=3(x+2)^2-4. Sure! Go for it! :)

OpenStudy (anonymous):

 i will substitute \[g(x)=3x ^{2}-4?\]

OpenStudy (joannablackwelder):

g(x)=3x^2-4 Yes. Now what does g(x^2)=?

OpenStudy (anonymous):

i dont understand how :/